OpenStudy (anonymous):

Consider the following expression: m^3+3m^2-2m+24 ____________________ m^2-5+5 Divide the polynomial in the numerator by the trinomial in the denomiator

OpenStudy (anonymous):

\[m^3+3m^2-2m+24\]---------------------- =\[m^2-5+5 = m^2\] \[m^3/m^2 + 3m^2/m^2-2m/m^2+24/m^2\] = m + 3 - 2/m + 24/m^2